# Revisiting wideband pulsar timing measurements

### v2

@@ 4 Demonstration using InPTA observations of PSR J2124–3358
We now proceed to demonstrate our new wideband measurement method using frequency-resolved integrated pulse profiles of PSR J2124–3338 and compare it against the wideband measurement method of PennucciDemorestRansom2014. PSR J2124–3358 is a bright, nearby, and isolated millisecond pulsar with a rotational period $P_{s}\sim$ 4.93 ms, located at a distance of around 400 pc. It was discovered using the Parkes 64-m Radio Telescope in the Parkes 436 MHz survey of the southern sky (BailesJohnston+1997). Its stable, precisely timed pulses make it an important target for pulsar timing array (PTA) experiments (e.g. RanaTarafdar+2025).
The data used in this section include integrated pulse profiles obtained from observations done using the upgraded Giant Metre-wave Radio Telescope (uGMRT: GuptaAjithkumar+2017) as part of the InPTA experiment (JoshiArumugasamy+2018) during observation cycles 37–48 (MJD 58781–60896) in the 300–500 MHz band (Band 3). TOAs derived from subsets of this data were included in the first and second data releases of the InPTA (TarafdarNobleson+2022; RanaTarafdar+2025); however, the epochs 60426–60896 have not been published elsewhere. The observation and data reduction procedures used to obtain this dataset are described in SusobhananMaan+2021, TarafdarNobleson+2022, and RanaTarafdar+2025. We have excluded observing epochs significantly affected by RFI from this analysis.
The wideband measurement technique described in PennucciDemorestRansom2014 and implemented in the PulsePortraiture package is henceforth referred to as the standard timing technique. In the standard technique, the noise parameters $\sigma_{\alpha}$ are estimated by considering the mean of the last 25% [^3] This fraction (25%) can be adjusted while running PulsePortraiture. Unfortunately, there is no rigorous procedure to determine the optimal value of this fraction, so we have persisted with the default value used in PulsePortraiture. Further, this value has been used in previous InPTA wideband analyses (NoblesonAgarwal+2022; TarafdarNobleson+2022). of the Fourier bins in the power spectrum computed from $\tilde{P}_{\alpha k}$. We have implemented the amplitude and noise-marginalized likelihood given in equation (10) in PulsePortraiture, and this is referred to as MLAN [^4] Marginalized Likelihood over Amplitude and Noise hereafter.
@@
We performed timing analysis on the two sets of TOAs using the PINT package (LuoRansom+2021; SusobhananKaplan+2024). We start our initial timing analysis using the pulsar ephemeris published in RanaTarafdar+2025. The timing model includes pulsar spin-down parameterized by the frequency F0 and the frequency derivative F1, secular DM evolution parameterized by an average DM and two derivatives (DM1, DM2), and astrometric delays with the sky location (RA, DEC) as free parameters. These are the same parameters that are fitted in the InPTA Data Release 2 narrowband dataset (RanaTarafdar+2025). Proper motion and parallax are not fitted, but fixed to values taken from the European Pulsar Timing Array Data Release 2 (AntoniadisBabak+2023). The TOA and DM residuals for standard and MLAN methods are plotted in Figure 6.
@@
After the initial timing, we performed single-pulsar noise analysis (SPNA) using a linearized timing model following the method described in SusobhananVan_Haasteren2025. We recall that the Bayes’ theorem for SPNA is given by
@@
We implemented the posterior distribution given in equation (13) with the help of PINT. We model inaccuracies in TOA measurement uncertainties and pulse jitter using an EFAC (error factor) and an EQUAD (error added in quadrature) (KikunagaHisano+2024). Similarly, inaccuracies in DM measurement uncertainties are modeled using a DMEFAC. We model stochastic DM variations and the achromatic red noise (due to rotational irregularities and gravitational wave background; AgazieAntoniadis+2024) as Fourier series Gaussian processes with power-law spectra. We use 50 linearly spaced frequency bins each for these Gaussian process models with a fundamental frequency $f_{1}=T_{\text{span}}^{-1}$ where $T_{\text{span}}=2087.5$ days. Following Van_HaasterenVallisneri2014b, we also include four logarithmically spaced frequency bins below $f_{1}$ to better capture the lower frequency components of the spin noise and DM noise (at $f_{1}/2$, $f_{1}/4$, $f_{1}/8$, and $f_{1}/16$). We perform parameter estimation on the wideband measurements using the above-mentioned model with the help of the nautilus package, which implements a neural network-boosted importance nested sampling algorithm (nautilus). We perform this analysis on wideband measurements obtained using both the standard and MLAN methods.
@@ 5 Summary and Discussion
We derived a new method for performing wideband TOA and DM measurements from frequency-resolved integrated pulse profiles of a pulsar following a Bayesian approach. Our method, compared to the widely used wideband measurement method of PennucciDemorestRansom2014, features a more rigorous treatment of noise in each frequency channel of the profile, leading to more realistic TOA and DM measurement uncertainty estimates. Our method, similar to PennucciDemorestRansom2014, allows one to make independent TOA and DM measurements by choosing an appropriate fiducial observing frequency. We demonstrated our method using a simulated profile and showed that (a) it accurately recovers the injected parameter values, and (b) the fiducial observing frequency selection procedure is effective in removing the TOA-DM measurement covariance. We applied our method to the InPTA data of PSR J2124–3318, and compared its performance against the method of PennucciDemorestRansom2014. We showed that the TOA and DM measurement uncertainty estimates obtained using our method are indeed more realistic than those obtained using PennucciDemorestRansom2014. This is crucial for achieving a robust detection of nanohertz gravitational waves using the wideband timing method, since imprecise white noise estimates can lead to biases in pulsar timing array inference (AntoniadisArumugam+2024).
In the coming years and decades, as the PTA datasets grow, the wideband technique will be crucial for ensuring that the computational cost of performing PTA analysis remains manageable. This work is a crucial stepping stone towards ensuring the reliability of the wideband timing method, so that it can be applied to high-precision timing experiments like pulsar timing arrays. We plan to apply our method to the full InPTA dataset in the near future as part of the second InPTA wideband data release.
@@ Acknowledgments
We thank the staff of the GMRT who made our observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. SJD is supported by IBS under the project code IBS-R018-D1. SD is supported by SERB grant MTR/2023/000384. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886 and 1002/2/2023/RRC/R&D-II/14369). KR is supported by UGC JRF fellowship. AdS is supported by UGC JRF fellowship HT is supported by DST INSPIRE Fellowship, INSPIRE code IF210656. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, and 24H01813, and Bilateral Joint Research Projects of JSPS (120237710). KV is supported by CSIR JRF Fellowship. ZZ is supported by the Prime Minister’s Research Fellows scheme, Ref. No. TF/PMRF-22-7307.
@@
PulsePortraiture (PennucciDemorestRansom2014; Pennucci2019), PINT LuoRansom+2021; SusobhananKaplan+2024, nestle (Barbary2021), nautilus (nautilus), corner (Foreman-Mackey2016), numpy (HarrisMillman+2020), scipy (VirtanenGommers+2020), matplotlib (Hunter2007)
@@ Appendix B Derivation of the amplitude and noise-maximized likelihood for wideband measurements
Following PennucciDemorestRansom2014, we maximize equation (6) with respect to the $a_{\alpha}$ parameters, and and this yields

### v3

@@ 4 Demonstration using InPTA observations of PSR J2124–3358
We now proceed to demonstrate our new wideband measurement method using frequency-resolved integrated pulse profiles of PSR J2124–3338 and compare it against the wideband measurement method of Pennucci et al. (2014). PSR J2124–3358 is a bright, nearby, and isolated millisecond pulsar with a rotational period $P_{s}\sim$ 4.93 ms, located at a distance of around 400 pc. It was discovered using the Parkes 64-m Radio Telescope in the Parkes 436 MHz survey of the southern sky (Bailes et al., 1997). Its stable, precisely timed pulses make it an important target for pulsar timing array (PTA) experiments (e.g. Rana et al., 2025).
The data used in this section include integrated pulse profiles obtained from observations done using the upgraded Giant Metre-wave Radio Telescope (uGMRT: Gupta et al., 2017) as part of the InPTA experiment (Joshi et al., 2018) during observation cycles 37–48 (MJD 58781–60896) in the 300–500 MHz band (Band 3). TOAs derived from subsets of this data were included in the first and second data releases of the InPTA (Tarafdar et al., 2022; Rana et al., 2025); however, the epochs 60426–60896 have not been published elsewhere. The observation and data reduction procedures used to obtain this dataset are described in Susobhanan et al. (2021), Tarafdar et al. (2022), and Rana et al. (2025). We have excluded observing epochs significantly affected by RFI from this analysis.
The wideband measurement technique described in Pennucci et al. (2014) and implemented in the PulsePortraiture package is henceforth referred to as the standard timing technique. In the standard technique, the noise parameters $\sigma_{\alpha}$ are estimated by considering the mean of the last 25% [^4] This fraction (25%) can be adjusted while running PulsePortraiture. Unfortunately, there is no rigorous procedure to determine the optimal value of this fraction, so we have persisted with the default value used in PulsePortraiture. Further, this value has been used in previous InPTA wideband analyses (Nobleson et al., 2022; Tarafdar et al., 2022). of the Fourier bins in the power spectrum computed from $\tilde{P}_{\alpha k}$. We have implemented the amplitude and noise-marginalized likelihood given in equation (11) in PulsePortraiture, and this is referred to as MLAN [^5] Marginalized Likelihood over Amplitude and Noise hereafter.
@@
We performed timing analysis on the two sets of TOAs using the PINT package (Luo et al., 2021; Susobhanan et al., 2024). We start our initial timing analysis using the pulsar ephemeris published in Rana et al. (2025). The timing model includes pulsar spin-down parameterized by the frequency F0 and the frequency derivative F1, secular DM evolution parameterized by an average DM and two derivatives (DM1, DM2), and astrometric delays with the sky location (RA, DEC) as free parameters. These are the same parameters that are fitted in the InPTA Data Release 2 narrowband dataset (Rana et al., 2025). Proper motion and parallax are not fitted, but fixed to values taken from the European Pulsar Timing Array Data Release 2 (Antoniadis et al., 2023). The TOA and DM residuals for standard and MLAN methods are plotted in Figure 6.
@@
After the initial timing, we performed single-pulsar noise analysis (SPNA) using a linearized timing model following the method described in Susobhanan and van Haasteren (2025). We recall that the Bayes’ theorem for SPNA is given by
@@
We implemented the posterior distribution given in equation (14) with the help of PINT. We model inaccuracies in TOA measurement uncertainties and pulse jitter using an EFAC (error factor) and an EQUAD (error added in quadrature) (Kikunaga et al., 2024). Similarly, inaccuracies in DM measurement uncertainties are modeled using a DMEFAC. We model stochastic DM variations and the achromatic red noise (due to rotational irregularities and gravitational wave background; Agazie et al. (2024)) as Fourier series Gaussian processes with power-law spectra. We use 50 linearly spaced frequency bins each for these Gaussian process models with a fundamental frequency $f_{1}=T_{\text{span}}^{-1}$ where $T_{\text{span}}=2087.5$ days. Following van Haasteren and Vallisneri (2014), we also include four logarithmically spaced frequency bins below $f_{1}$ to better capture the lower frequency components of the spin noise and DM noise (at $f_{1}/2$, $f_{1}/4$, $f_{1}/8$, and $f_{1}/16$). We perform parameter estimation on the wideband measurements using the above-mentioned model with the help of the nautilus package, which implements a neural network-boosted importance nested sampling algorithm (Lange, 2023). We perform this analysis on wideband measurements obtained using both the standard and MLAN methods.
@@
Figure 8: Deviation of timing model parameters estimated from the standard dataset and the MLAN dataset. The parameter estimates are obtained after applying the median noise parameters obtained from SPNA.
@@ 5 Summary and Discussion
We derived a new method for performing wideband TOA and DM measurements from frequency-resolved integrated pulse profiles of a pulsar following a Bayesian approach. Our method, compared to the widely used wideband measurement method of Pennucci et al. (2014), features a more rigorous treatment of noise in each frequency channel of the profile, leading to more realistic TOA and DM measurement uncertainty estimates. Our method, similar to Pennucci et al. (2014), allows one to make independent TOA and DM measurements by choosing an appropriate fiducial observing frequency. We demonstrated our method using a simulated profile and showed that (a) it accurately recovers the injected parameter values, and (b) the fiducial observing frequency selection procedure is effective in removing the TOA-DM measurement covariance. We applied our method to the InPTA data of PSR J2124–3318, and compared its performance against the method of Pennucci et al. (2014). We showed that the TOA and DM measurement uncertainty estimates obtained using our method are indeed more realistic than those obtained using Pennucci et al. (2014). This is crucial for achieving a robust detection of nanohertz gravitational waves using the wideband timing method, since imprecise white noise estimates can lead to biases in pulsar timing array inference (Antoniadis et al., 2024).
In the coming years and decades, as the PTA datasets grow, the wideband technique will be crucial for ensuring that the computational cost of performing PTA analysis remains manageable. This work is a crucial stepping stone towards ensuring the reliability of the wideband timing method, so that it can be applied to high-precision timing experiments like pulsar timing arrays. We plan to apply our method to the full InPTA dataset in the near future as part of the second InPTA wideband data release. However, we note that several major challenges remain, including (1) the rigorous handling of pulse broadening due to interstellar scattering (Sutton, 1971), and (2) the handling of mode changes (Backer, 1970) and other types of profile variability events (e.g. Singha et al., 2021). In addition, we found a few epochs that were severely affected by interstellar scintillation, such that there was very low S/N in some parts of the band, but very high S/N in others. In such epochs, we found that both the MLAN and standard methods produced imperfect fits to the observed portrait, and the TOAs and DMs estimated by the two methods showed significant differences. We are currently working on tackling these challenges.
@@ Acknowledgments
We thank the staff of the GMRT who made our observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. SJD is supported by IBS under the project code IBS-R018-D1. SD is supported by SERB grant MTR/2023/000384. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886 and 1002/2/2023/RRC/R&D-II/14369). KR is supported by UGC JRF fellowship. AdS is supported by UGC JRF fellowship HT is supported by DST INSPIRE Fellowship, INSPIRE code IF210656. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, and 24H01813, and Bilateral Joint Research Projects of JSPS (120237710). KV is supported by CSIR JRF Fellowship. ZZ is supported by the Prime Minister’s Research Fellows scheme, Ref. No. TF/PMRF-22-7307. AS thanks Rutger van Haasteren for fruitful discussions.
@@
PulsePortraiture (Pennucci et al., 2014; Pennucci, 2019), PINT (Luo et al., 2021; Susobhanan et al., 2024), nestle (Barbary, 2021), nautilus (Lange, 2023), corner (Foreman-Mackey, 2016), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007)
@@ Appendix B Derivation of the amplitude and noise-maximized likelihood for wideband measurements
Following Pennucci et al. (2014), we maximize equation (6) with respect to the $a_{\alpha}$ parameters, and and this yields